Does the line below pass through the point (2, -1)?
3x – 2y = 8

Answer by funmath(2933)

Does the line below pass through the point (2, -1)?
3x – 2y = 8
:
The ordered pair (x,y)=(2,-1). Substitute 2 for x and -1 for y and see the both sides equal. If it does, then the point is on the the line. If not, it's not.
3(2)-2(-1)=8 ?
6+2=8
8=8 !
Yes (2,-1) is on the line.
Happy Calculating!!!